Please solve this question, 
A cash box contains some coins to the value of £5.25.
There are twice as many 5p coins as 2p coins and twice as many 2p coins as 1p coins.
This means there are:
------- 5p coins ---------2p -------1p coins 
Regards

Found 3 solutions by rothauserc, MathTherapy, greenestamps:
Answer by rothauserc(4718)

let x be the number of 1p coins
:
number of 2p coins is 2x
:
number of 5p coins is 4x
:
100 pence makes one pound(as of 1971)
:
x +2(2x) +5(4x) = 525 pence
:
25x = 525
:
x = 21
:
******************************************
84 5p coins, 42 2p coins, 21 1p coins
******************************************
:

Answer by MathTherapy(10559)


Please solve this question, 
A cash box contains some coins to the value of £5.25.
There are twice as many 5p coins as 2p coins and twice as many 2p coins as 1p coins.
This means there are:
------- 5p coins ---------2p -------1p coins 
Regards
Correct answer: 
ALL other answers are INCORRECT, so IGNORE them!

Answer by greenestamps(13216)



Solving this problem using formal algebra is of course a good exercise in learning the methods of algebra. But solving it informally using logical reasoning and a bit of mental arithmetic can also be good exercise for the brain.

The number of 2p coins is twice the number of 1p coins, and the number of 5p coins is twice the number of 2p coins. So for each 1p coin there are 2 2p coins and 4 5p coins.

So picture groups of coins consisting of 4 5p coins, 2 2p coins, and 1 1p coin. The value of such a group of coins is 4(5)+2(2)+1(1) = 25p.

The number of those groups required to make a total of £5.25 = 525p is 525/25 = 21.

21 of the groups means 21 1p coins, 42 2p coins, and 42 5p coins.